# The 2D free particle in the phase space quantum mechanics

### v3

@@ Abstract
The Wigner eigenfunctions of a free quantum particle propagating on a plane are derived. Two possibilities are analysed. First, the particle with fixed components of the Cartesian momentum is considered. Then the particle of given energy and angular momentum is discussed. In that second case, a special choice of coordinates on the symplectic space $(\mathbb{R}^{4},\,\omega)$ suitable for the representation of eigenstates of the discussed particle is presented. Further, the Moyal $\star_{(\text{M})}$-product on the phase space is derived with the use of the Fedosov algorithm adapted to these coordinates on a flat phase space. Next, the eigenvalue equations for the Wigner eigenfunction are solved and the physically acceptable solutions are identified. Finally, a relationship between the Wigner eigenfunction of the particle with the fixed components of the Cartesian momentum and the cross-Wigner functions of the particle with the given energy and angular momentum is found.
@@ 1 Introduction
In the current article, we extend that approach to solve the star-eigenvalue problem for a free quantum particle of mass $M$ moving in the space $\mathbb{R}^{2}$. The particle is characterised either by components of the Cartesian momentum or by energy and the angular momentum.
The first option is easily solvable. The second case is much more demanding. In order to deal with it we choose some special Darboux coordinates $(T,\chi,H,L)$. Then employing the adapted Fedosov algorithm, we explicitly construct the $\star_{(\text{M})}$-product in the new chart. We formulate the $\star_{(\text{M})}$-eigenvalue problem for both the Hamiltonian $H$ and the angular momentum $L$ and solve two $\star_{(\text{M})}$-eigenvalue problems explicitly, obtaining a collection of Wigner eigenfunctions $W_{Em}.$ Finally, we examine which of the resulting functions represent physically acceptable Wigner eigenfunctions of the 2D free particle of given energy $E$ and the magnetic quantum number $m$.
The last part of our work refers to the vital question of how to decompose states in combinations of some quasiprobability distributions. Namely, knowing the Wigner eigenfunction $W_{p_{x0}p_{y0}}$ for the 2D free molecule with the fixed components of the Cartesian momentum $p_{x0},p_{y0}$ and some cross-Wigner functions $W_{Emm^{\prime}}$ for the 2D free particle with the defined energy $E$ and angular momenta $m\hbar,m^{\prime}\hbar$ we find an expansion of $W_{p_{x0}p_{y0}}$ in the generalised linear combination of functions from the set $\{W_{Emm^{\prime}}\}_{E>0,m,\,m^{\prime}\in\mathbb{Z}}.$ This result illustrates the structure of set of states as a linear space.
2 Wigner eigenfunctions of Cartesian momentum and energy–angular momentum
@@
Eigenstates of a free quantum particle living on the plane $\mathbb{R}^{2}$ are characterised either by values of components of the momentum $p_{x},p_{y}$ or the energy being value of the Hamilton function $H$ and the angular momentum $L.$ These two kinds of eigenstates are not equivalent i.e. a quantum particle with defined components of the momentum has an undetermined angular momentum and vice versa.
Let us discuss the aforementioned eigenstates of the 2D free particle of mass $M$ in frames of phase space quantum mechanics. The quantum phase space of this particle is—like in the classical physics—the symplectic manifold $(\mathbb{R}^{4},\,\omega).$
We need to introduce the Moyal product on the symplectic manifold. This product is a counterpart of composition of linear operators in a Hilbert space. The Moyal product $\star_{(\text{M})}$ of functions $f(\mathbf{r},\,\mathbf{p})$ and $g(\mathbf{r},\,\mathbf{p})$ is defined (compare [18, 17]) in the Cartesian chart $(\mathbf{r},\,\mathbf{p})\equiv(x,\,y,\,p_{x},\,p_{y})$ as
@@
In several cases however, one can apply the differential formula
@@
As we already mentioned, eigenstates of quantum 2D free particle can be determined by two complete sets of observables $\{R_{1},\,R_{2}\}$ commuting in the sense of the Moyal bracket:
@@
    the components of the momentum: $R_{1}:=p_{x}$ and $R_{2}:=p_{y}$,
@@
    the energy and the angular momentum: $R_{1}:=H=\frac{p^{2}_{x}+p^{2}_{y}}{2M}$ and $R_{2}:=L=xp_{y}-yp_{x}$.
@@
$\{p_{x},\,L\}_{(\text{M})}\neq 0\;\;\;,\;\;\;\{p_{y},\,L\}_{(\text{M})}\neq 0.$
The Wigner eigenfunction $W_{r_{1}r_{2}}(\mathbf{r},\,\mathbf{p})$ of the 2D free particle is a real function or a tempered distribution on $\mathbb{R}^{4}$ that satisfies the system of eigenvalue equations
@@
$\{R_{2},\,W_{r_{1}r_{2}}\}_{(\text{M})}(\mathbf{r},\,\mathbf{p})$ = 0. (2.5d)
By $r_{1}$ and $r_{2}$ we denote the eigenvalues of functions $R_{1}$ and $R_{2}$ respectively. Moreover, the Wigner eigenfunctions $W_{r_{1}r_{2}}(\mathbf{r},\,\mathbf{p})$ have to obey the condition of positivity of the marginal distributions
$\int_{\mathbb{R}^{2}}\text{d}\mathbf{r}W_{r_{1}r_{2}}(\mathbf{r},\,\mathbf{p})$ $\geq 0$ , $\int_{\mathbb{R}^{2}}\text{d}\mathbf{p}W_{r_{1}r_{2}}(\mathbf{r},\,\mathbf{p})$ $\geq 0$ (2.6)
and some normalization condition which explicit form will be derived in the next section
@@
In the case when components of the momentum of particle are determined, the solution of Eqs. (2.5a)–(2.5d) is well known
Symbols $p_{x0},\,p_{y0}\in\mathbb{R}$ denote eigenvalues of the Cartesian components of the momentum. One can check that the Wigner eigenfunction (2.9) also satisfies conditions (2.6).
Let us consider the second option—the 2D particle with the defined energy and angular momentum. One can see that $R_{1}=H$ and $R_{2}=L$ are quadratic functions of position and momentum. Hence, the Moyal brackets in Eqs. (2.5c)–(2.5d) are proportional to the respective Poisson brackets. This means that if we find coordinates $T$ and $\chi$ such that the transformation $(x,\,y,\,p_{x},p_{y})\to(T,\,\chi,\,H,\,L)$ is symplectic, then the LHS of Eqs. (2.5c)–(2.5d) reveals its simplest form: $\{\cdot,\,H\}_{(\text{M})}=\frac{\partial}{\partial T}(\cdot)$ and $\{\cdot,\,L\}_{(\text{M})}=\frac{\partial}{\partial\chi}(\cdot)$. In order to derive the explicit expressions of $T$ and $\chi$, one considers conservation of the symplectic form $\omega$ under the aforementioned transformation. One obtains that
@@
The transformation (2.10) is well-defined only if $p_{x}^{2}+p_{y}^{2}>0$ i.e. it does not work for the classical particles at rest. Fortunately, the subset $p_{x}^{2}+p^{2}_{y}=0$ is a set of measure zero. Hence, it can be ignored from a quantum point of view. One can check that the Jacobian of the transformation equals
@@
Assuming $p_{x}^{2}+p_{y}^{2}>0$, the inverse transformation is given
@@
Notice that the function $W_{Em}$ written in the form (2.23) depends exclusively on the variables $H$ and $L$ representing energy and classical angular momentum.
Lets prove that the solution (2.23) is the Wigner eigenfunction only if $m\in\mathbb{Z}$. In order to do that we use the coordinates $(r,\,\phi,\,p,\,\chi)$, where the pair $(r,\,\phi)$ denotes the polar coordinates of the position vector $\mathbf{r}$ and $(p,\,\chi)$ are the polar coordinates of the momentum vector $\mathbf{p}$. These new coordinates are not symplectic. The elementary volume $\text{d}T\text{d}\chi\text{d}H\text{d}L$ transforms into $r\text{d}r\text{d}\phi p\text{d}p\text{d}\chi.$
The energy is of the form $E=\frac{p^{2}_{0}}{2M}$, where $p_{0}$ denotes the length of some momentum vector. Then the function (2.23) turns into
@@
By $J_{n}(x),\;n\in\mathbb{Z}$ we mean the Bessel function of the first kind. One can see that function $P_{Em}$ does not depend on $\phi$. We introduce $\theta$ fulfilling the conditions
@@
Considering the condition (2.30) and the result of Appendix B i.e.
$\big(W_{Em}\star_{(\text{M})}W_{\widetilde{E}m}\big)(T,\,\chi,\,H,\,L)=\pi N_{Em}\delta(E-\widetilde{E})W_{Em}(T,\,\chi,\,H,\,L),$ (2.31)
we obtain
@@
Let us assume $m\in\mathbb{R}\setminus\mathbb{Z}$. Unfortunately, in that case, we do not know the closed form of the integral (2.28). We choose the same $N_{Em}$ as (2.32) and do numerical calculations. The figure 1 illustrates that the integral (2.28) for a non-integer quantum angular momentum number $m$ becomes negative for some radii $r$. It contradicts the interpretation of $P_{Em}(r,\,\phi)$ as a spatial density of probability.
@@
In the previous chapter, we have derived the Wigner eigenfunction $W_{Em}(\mathbf{r},\,\mathbf{p})$, given by Eq. (2.23). In fact, it represents the state $|\Psi_{Em}\rangle$ with a given energy $E>0$ and angular momentum $m\hbar$, where $m\in\mathbb{Z}$. Analogously, the Wigner eigenfunction $W_{p_{x0}p_{y0}}(\mathbf{r},\,\mathbf{p})$ (see Eq. (2.9)) represents the state $|\Psi_{p_{x0}p_{y0}}\rangle$ with a given Cartesian momentum $(p_{x0},\,p_{y0})\in\mathbb{R}^{2}$. We will denote the cross Wigner function of the above states $|\Psi_{Em}\rangle$ and $|\Psi_{Em^{\prime}}\rangle$ as $W_{Emm^{\prime}}(\mathbf{r},\,\mathbf{p})$. It is defined for each $E>0,\,m,m^{\prime}\in\mathbb{Z}$ as a solution of the system of star-eigenvalue equations
@@
In the coordinates $(T,\,\chi,\,H,\,L)$ defined by the relations (2.10), the formulas (3.1a) and (3.1d) look like Eqs. (2.16a) and (2.16c). Therefore we postulate that
@@
In Appendix B we have proved that
$\big(W_{Emm^{\prime}}\star_{(\text{M})}W_{\widetilde{E}m^{\prime}\widetilde{m}^{\prime}}\big)(T,\,\chi,\,H,\,L)=\frac{\pi N_{Emm^{\prime}}N_{\widetilde{E}m^{\prime}\widetilde{m}^{\prime}}}{N_{Em\widetilde{m}^{\prime}}}\delta(E-\widetilde{E})W_{Em\widetilde{m}^{\prime}}(T,\,\chi,\,H,\,L).$ (3.7)
@@
Using the property $|z^{n}|=|z|^{n}$ for $n\in\mathbb{Z}$, Eq. (3.9) induces that $|z|=1$. Therefore, the normalization constant can be determined up to a phase factor $\alpha\in\mathbb{R}$
@@
$W_{\widetilde{p}_{x0}\widetilde{p}_{y0}}(T,\,\chi,\,H,\,L)=\sum^{\infty}_{\widetilde{m}=-\infty}\sum^{\infty}_{\widetilde{m}^{\prime}=-\infty}C_{\widetilde{m}\widetilde{m}^{\prime}}W_{\widetilde{E}\widetilde{m}\widetilde{m}^{\prime}}(T,\,\chi,\,H,\,L),$ (3.11)
where the Wigner eigenfunction $W_{\widetilde{p}_{x0}\widetilde{p}_{y0}}$ is defined as (2.9). In the coordinates (2.10), this function looks like
The coefficients in the formula (3.12) are equal to
@@
In Appendix C, we have shown that the LHS of Eq. (3.14) equals
@@
Statement
The work was partially financed under the program ‘FU2N - Fund for the Improvement of Skills of Young Scientists’ supporting the research excellence of employees and doctoral students of the Lodz University of Technology.

### v4

@@ Abstract
The Wigner eigenfunctions of a free quantum particle propagating on a plane are derived. Two possibilities are analysed. Firstly, the particle of given energy and angular momentum is discussed. In that case, a special choice of coordinates on the symplectic space $(\mathbb{R}^{4},\,\omega)$ suitable for the representation of eigenstates of the discussed particle is presented. Further, the Moyal $\star_{(\text{M})}$-product on the phase space is derived with the use of the Fedosov algorithm adapted to these coordinates on a flat phase space. Next, the eigenvalue equations for the Wigner eigenfunction are solved and the physically acceptable solutions are identified. Secondly, the particle with fixed components of the Cartesian momentum is considered. Finally, a relationship between the Wigner eigenfunction of the particle with the fixed components of the Cartesian momentum and the cross-Wigner functions of the particle with the given energy and angular momentum is found.
@@ 1 Introduction
In the current article, we extend that approach to solve the star-eigenvalue problem for a free quantum particle of mass $M$ moving in the space $\mathbb{R}^{2}$. The particle is characterised either by energy and the angular momentum or by components of the Cartesian momentum.
The second option is easily solvable. The first case is much more demanding. In order to deal with it we choose some special Darboux coordinates $(T,\chi,H,L)$. Then employing the adapted Fedosov algorithm, we explicitly construct the $\star_{(\text{M})}$-product in the new chart. We formulate the $\star_{(\text{M})}$-eigenvalue problem for both the Hamiltonian $H$ and the angular momentum $L$ and solve two $\star_{(\text{M})}$-eigenvalue problems explicitly, obtaining a collection of Wigner eigenfunctions $W_{Em}.$ Finally, we examine which of the resulting functions represent physically acceptable Wigner eigenfunctions of the 2D free particle of given energy $E$ and the magnetic quantum number $m$.
The last part of our work refers to the vital question of how to decompose states in combinations of some quasiprobability distributions. Namely, knowing the Wigner eigenfunction $W_{p_{x0}p_{y0}}$ for the 2D free molecule with the fixed components of the Cartesian momentum $p_{x0},p_{y0}$ and some cross-Wigner functions $W_{Emm^{\prime}}$ for the 2D free particle with the defined energy $E$ and angular momenta $m\hbar,m^{\prime}\hbar$ we find an expansion of $W_{p_{x0}p_{y0}}$ in the generalised linear combination of functions from the set $\{W_{Emm^{\prime}}\}_{E>0;m,\,m^{\prime}\in\mathbb{Z}}.$ This result illustrates the structure of set of states as a linear space.
2 The Wigner eigenfunctions of states with a fixed energy and an angular momentum and the Wigner eigenfunctions of states of a given Cartesian momentum
@@
States of a free quantum particle living on the plane $\mathbb{R}^{2}$ are characterised either by values of the energy being values of the Hamilton function $H$ and of the angular momentum $L$ or alternatively, by components of the Cartesian momentum $p_{x},p_{y}$. In contrast to classical mechanics, these two kinds of quantum states are not equivalent i.e. the quantum particle with defined components of the Cartesian momentum has an undetermined angular momentum and vice versa.
The quantum phase space of this particle is—like in the classical physics—the symplectic manifold $(\mathbb{R}^{4},\,\omega).$ We need to introduce the Moyal product on the symplectic manifold. This product is a counterpart of composition of linear operators in a Hilbert space. The Moyal product $\star_{(\text{M})}$ of functions $f(\mathbf{r},\,\mathbf{p})$ and $g(\mathbf{r},\,\mathbf{p})$ is defined (compare [19, 18]) in the Cartesian chart $(\mathbf{r},\,\mathbf{p})=(x,\,y,\,p_{x},\,p_{y})$ as
@@
However, in several cases one can apply the differential formula
@@
As we already mentioned, the eigenstates of quantum 2D free particle can be determined by two complete sets of observables $\{R_{1},\,R_{2}\}$ commuting in the sense of the Moyal bracket:
@@
    the energy and the angular momentum: $R_{1}:=H=\frac{p^{2}_{x}+p^{2}_{y}}{2M}$ and $R_{2}:=L=xp_{y}-yp_{x}$,
@@
    the components of the momentum: $R_{1}:=p_{x}$ and $R_{2}:=p_{y}$.
@@
$\{p_{x},\,L\}_{(\text{M})}\neq 0$ and $\{p_{y},\,L\}_{(\text{M})}\neq 0$.
The Wigner eigenfunction $W_{r_{1}r_{2}}(\mathbf{r},\,\mathbf{p})$ of the 2D free particle is a real function or a tempered distribution on $\mathbb{R}^{4}$ that satisfies the system of $\star_{(\text{M})}$-eigenvalue equations
@@
$\{R_{2},\,W_{r_{1}r_{2}}\}_{(\text{M})}(\mathbf{r},\,\mathbf{p})$ = 0 (2.5d)
and the normalization condition
@@
By $r_{1}$ and $r_{2}$ we denote the eigenvalues of functions $R_{1}$ and $R_{2}$ respectively.
Moreover, the Wigner eigenfunctions $W_{r_{1}r_{2}}(\mathbf{r},\,\mathbf{p})$ have to obey the condition of positivity of the marginal distributions
$\int_{\mathbb{R}^{2}}\text{d}\mathbf{r}W_{r_{1}r_{2}}(\mathbf{r},\,\mathbf{p})$ $\geq 0,$ $\int_{\mathbb{R}^{2}}\text{d}\mathbf{p}W_{r_{1}r_{2}}(\mathbf{r},\,\mathbf{p})$ $\geq 0.$ (2.8)
@@
Let us consider the first option—the 2D particle with the defined energy and angular momentum. One can see that $R_{1}=H$ and $R_{2}=L$ are quadratic functions of position and momentum. Hence, the Moyal brackets in Eqs. (2.5c)–(2.5d) are proportional to the respective Poisson brackets. This means that if we find coordinates $T$ and $\chi$ such that the transformation $(x,\,y,\,p_{x},p_{y})\to(T,\,\chi,\,H,\,L)$ is symplectic, then the LHS of Eqs. (2.5c)–(2.5d) reveals its simplest form: $\{\cdot,\,H\}_{(\text{M})}=\frac{\partial}{\partial T}(\cdot)$ and $\{\cdot,\,L\}_{(\text{M})}=\frac{\partial}{\partial\chi}(\cdot)$. In order to derive the explicit expressions of $T$ and $\chi$, one considers conservation of the symplectic form $\omega$ under the aforementioned transformation. One obtains that
@@
The transformation (2.9) is well-defined only if $p_{x}^{2}+p_{y}^{2}>0$ i.e. it does not work for the classical particles at rest. Fortunately, the subset $p_{x}^{2}+p^{2}_{y}=0$ is a set of measure zero. Hence, it can be ignored from the quantum point of view. One can check that the Jacobian of the transformation equals
@@
Assuming $p_{x}^{2}+p_{y}^{2}>0$, the inverse transformation is given by expressions
@@
The function $W_{Em}$ written in the form (2.22) was first derived in [15].
Lets prove that the solution (2.22) is the Wigner eigenfunction only if $m\in\mathbb{Z}$. We are going to do that in two ways: numerically and analytically.
We discuss the numerical method first. Let us introduce the coordinates $(r,\,\phi,\,p,\,\chi)$, where the pair $(r,\,\phi)$ denotes the polar coordinates of the position vector $\mathbf{r}$ and $(p,\,\chi)$ are the polar coordinates of the momentum vector $\mathbf{p}$. These new coordinates are not symplectic. The elementary volume $\text{d}T\text{d}\chi\text{d}H\text{d}L$ transforms into $r\text{d}r\text{d}\phi p\text{d}p\text{d}\chi$. The energy is of the form $E=\frac{p^{2}_{0}}{2M}$, where $p_{0}$ denotes the length of some momentum vector. Then the function (2.22) turns into
@@
By $J_{n}(x),\;n\in\mathbb{Z}$ we mean the Bessel function of the first kind. One can see that function $P_{Em}$ does not depend on $\phi$. We introduce a new variable $\theta$ fulfilling the conditions
@@
Applying the result (B.22) from Appendix B we obtain
@@
Let us assume $m\in\mathbb{R}\setminus\mathbb{Z}$. Unfortunately, in that case, we do not know the closed form of the integral (2.27). We choose the value of the coefficient $N_{Em}$ as (2.30) and perform numerical calculations. At Fig.1 one can see that the integral (2.27) for a non-integer quantum angular momentum number $m$ becomes negative for some radii $r$. This observation contradicts the interpretation of $P_{Em}(r,\,\phi)$ as a spatial density of probability.
@@
The analytical method is based on considerations presented in Appendix B. We prove that only for $m,\,\widetilde{m}\in\mathbb{Z}$ the normalization condition (B.22) holds and the normalization coefficient equals (2.30).
@@
In the second case when components of the momentum of particle are determined, the solution of Eqs. (2.5a)–(2.5d) is well known
Symbols $p_{x0},\,p_{y0}\in\mathbb{R}$ denote eigenvalues of the Cartesian components of the momentum. In the coordinates (2.9), this function looks like
The parametres in the formula (2.32) are equal to
3 The cross-Wigner functions of energy-angular momentum
In the previous chapter, we have derived the Wigner eigenfunction $W_{Em}(\mathbf{r},\,\mathbf{p})$, given by Eq. (2.22). In fact, it represents the state $|\Psi_{Em}\rangle$ with a given energy $E>0$ and angular momentum $m\hbar$, where $m\in\mathbb{Z}$. Analogously, the Wigner eigenfunction $W_{p_{x0}p_{y0}}(\mathbf{r},\,\mathbf{p})$ (see Eq. (2.31)) represents the state $|\Psi_{p_{x0}p_{y0}}\rangle$ with a given Cartesian momentum $(p_{x0},\,p_{y0})\in\mathbb{R}^{2}$.
We will denote the cross-Wigner function of the above states $|\Psi_{Em}\rangle$ and $|\Psi_{Em^{\prime}}\rangle$ as $W_{Emm^{\prime}}(\mathbf{r},\,\mathbf{p})$. It is defined for each $E>0$ and $m,m^{\prime}\in\mathbb{R}$ as a solution of the system of star-eigenvalue equations
@@ 3 The cross-Wigner functions of energy-angular momentum
In the coordinates $(T,\,\chi,\,H,\,L)$ defined by the relations (2.9), the formulas (3.34a) and (3.34d) look like Eqs. (2.15a) and (2.15c) respectively. Therefore we postulate that
@@
In the Appendix B, we show that only if $m,\,m^{\prime},\,\widetilde{m},\,\widetilde{m}^{\prime}\in\mathbb{Z}$, the normalization condition holds (compare with Eq. (B.21))
$\big(W_{Emm^{\prime}}\star_{(\text{M})}W_{\widetilde{E}\widetilde{m}\widetilde{m}^{\prime}}\big)(T,\,\chi,\,H,\,L)=\frac{\pi N_{Emm^{\prime}}N_{\widetilde{E}\widetilde{m}\widetilde{m}^{\prime}}}{N_{Em\widetilde{m}^{\prime}}}\delta(E-\widetilde{E})W_{Em\widetilde{m}^{\prime}}(T,\,\chi,\,H,\,L)\delta_{m^{\prime}\widetilde{m}}$. (3.39)
@@
Using the property $|z^{n}|=|z|^{n}$ for $n\in\mathbb{Z}$, Eq. (3.42) induces that $|z|=1$. Therefore, the normalization constant can be determined up to a phase factor $\alpha\in\mathbb{R}$ as
@@
$W_{\widetilde{p}_{x0}\widetilde{p}_{y0}}(T,\,\chi,\,H,\,L)=\sum^{\infty}_{\widetilde{m}=-\infty}\sum^{\infty}_{\widetilde{m}^{\prime}=-\infty}C_{\widetilde{m}\widetilde{m}^{\prime}}W_{\widetilde{E}\widetilde{m}\widetilde{m}^{\prime}}(T,\,\chi,\,H,\,L).$ (4.1)
@@
In Appendix C, we prove that the LHS of Eq. (4.2) equals
@@
5 Conclusions
All considerations contained in the present paper are confined to the phase space quantum mechanics formalism exclusively. We derive the energy-angular momentum Wigner eigenfunction $W_{Em}$ and the cross-Wigner function $W_{Emm^{\prime}}$ by solving the $\star_{\text{(M)}}$-eigenvalue equations. We show that not every solution obtained this way is physically acceptable, but only those with the quantized magnetic number and satisfying the normalization condition.
It is worth mentioning that we construct the Moyal $\star_{(\text{M})}$-product with use of the Fedosov algorithm adapted to the symplectic coordinates $(T,\,\chi,\,H,\,L)$. We prove that the coordinates are the most suitable for energy-angular momentum $\star_{\text{M}}$-eigenproblem, because equations $\{H,\,W_{Em}\}_{(\text{M})}=0$ and $\{L,\,W_{Em}\}_{\text{(M)}}=0$ are satisfied by functions of the form $W_{Em}(T,\,\chi,\,H,\,L)=W_{Em}(H,\,L)$. It results from the fact that the Moyal bracket reduces to the Poisson bracket.
Finally, inspired by the Jacobi-Anger expansion in the Hilbert quantum mechanics, we expand the Wigner eigenfunction of Cartesian momentum $W_{\widetilde{p}_{x0}\widetilde{p}_{y0}}$ in the cross-Wigner functions of energy-angular momentum $W_{Emm^{\prime}}$. We achieve that by calculating the products $W_{Emm^{\prime}}\star_{\text{(M)}}W_{E\widetilde{m}\widetilde{m}^{\prime}}$ and $W_{Emm^{\prime}}\star_{\text{(M)}}W_{\widetilde{p}_{x0}\widetilde{p}_{y0}}$. An open question remains whether the proposed algorithm can be applied to other expansions known in the Hilbert space formalism, e.g. the interbasis expansions of the 2D hydrogen atom [22].